OpenStudy (anonymous):

2/3 cup of bleached flour and 1/4 cup of wheat flour. how many cups of bleached flour ?

OpenStudy (anonymous):

use 12 as your lcm

OpenStudy (anonymous):

8/12+3/12=11/12

OpenStudy (curry):

=11/12